OpenStudy (3psilon):

Physics help? No one seems to be in the physics section. 1. A cat jumps off of the table with a velocity v0 at an angle θ. The table is h meters tall. Write the velocity vector as a function of time. v(t)=? 2. What is the angle with respect to the ground that the cat makes when it hits the ground?

OpenStudy (3psilon):

What I got for number 1 \[\frac{ -v _{0}\sin \Theta \pm \sqrt{v _{0}\sin \Theta ^{2} + 19.6h} }{ -9.8 }\]

OpenStudy (3psilon):

I don't exactly understand number 2